What is kinetic enrgy of system which is in polar coordinates system.
1
Expert's answer
2020-12-06T17:30:03-0500

Answer

kinetic enrgy of system which is in polar coordinates system (r, "\theta")

Can be written as

"T=\frac{m}{2}(\dot{r}^2+r^2\dot{\theta}^2)"

Where

m is mass of body.